1a) what is the value of 10 Year K1000 par value bond with a 10% annual coupon, paid semi annually, if its required return is  10%

b) what is the value of a 13% coupon bond that is otherwise identical to the bond described in part c above? would we now have a discount or a premium?

c) what is the value of a 7% coupon with these characteristics ? would we now have a discount or premium bond?

d) what would happen to the 7%, 10% and 13% coupon bonds overtime if the required has increased to 13%
Answer #1

The formula for Value of bond:

,
where MP = market price
CF = cash flows
r = required rate of return
t = time period

a. Bond with 10% coupon

CF = cash flows = (10%*1000)/2 = 50
r = required rate of return = 10%/2 = 5%
t = time period = 1,2,....20

Time CF (1+r)^n CF/(1+r)^n Price = Sum of CF/(1+r)^n
1 50 1.05 47.6190476 1000
2 50 1.1025 45.3514739
3 50 1.157625 43.1918799
4 50 1.21550625 41.1351237
5 50 1.27628156 39.1763083
6 50 1.34009564 37.3107698
7 50 1.40710042 35.5340665
8 50 1.47745544 33.8419681
9 50 1.55132822 32.2304458
10 50 1.62889463 30.6956627
11 50 1.71033936 29.2339645
12 50 1.79585633 27.8418709
13 50 1.88564914 26.5160675
14 50 1.9799316 25.2533976
15 50 2.07892818 24.0508549
16 50 2.18287459 22.9055761
17 50 2.29201832 21.8148344
18 50 2.40661923 20.7760327
19 50 2.5269502 19.7866979
20 1050 2.65329771 395.733957

The above bond is selling at par.

b. Bond with 13% coupon

CF = cash flows = (13%*1000)/2 = 65
r = required rate of return = 10%/2 =5%
t = time period = 1,2,....20

Time CF (1+r)^n CF/(1+r)^n Price = Sum of CF/(1+r)^n
1 65 1.05 61.9047619 1186.933155
2 65 1.1025 58.9569161
3 65 1.157625 56.1494439
4 65 1.21550625 53.4756609
5 65 1.27628156 50.9292008
6 65 1.34009564 48.5040008
7 65 1.40710042 46.1942865
8 65 1.47745544 43.9945585
9 65 1.55132822 41.8995796
10 65 1.62889463 39.9043615
11 65 1.71033936 38.0041538
12 65 1.79585633 36.1944322
13 65 1.88564914 34.4708878
14 65 1.9799316 32.8294169
15 65 2.07892818 31.2661114
16 65 2.18287459 29.7772489
17 65 2.29201832 28.3592847
18 65 2.40661923 27.0088426
19 65 2.5269502 25.7227072
20 1065 2.65329771 401.387299

We would have a premium since the coupon rate is more than the required rate of return.

c. Bond with 7% coupon

CF = cash flows = (7%*1000)/2 = 35
r = required rate of return = 10%.2= 5%
t = time period = 1,2,....20

Time CF (1+r)^n CF/(1+r)^n Price = Sum of CF/(1+r)^n
1 35 1.05 33.3333333 813.0668449
2 35 1.1025 31.7460317
3 35 1.157625 30.2343159
4 35 1.21550625 28.7945866
5 35 1.27628156 27.4234158
6 35 1.34009564 26.1175389
7 35 1.40710042 24.8738466
8 35 1.47745544 23.6893777
9 35 1.55132822 22.5613121
10 35 1.62889463 21.4869639
11 35 1.71033936 20.4637751
12 35 1.79585633 19.4893096
13 35 1.88564914 18.5612473
14 35 1.9799316 17.6773784
15 35 2.07892818 16.8355984
16 35 2.18287459 16.0339033
17 35 2.29201832 15.2703841
18 35 2.40661923 14.5432229
19 35 2.5269502 13.8506885
20 1035 2.65329771 390.080615

The bond would sell at a discount as the coupon rate is less than the required rate of return.

d. If the rate of interest increases to 13% over time, the prices of the three bonds would reduce over time. But the reinvestment value of the coupon payments shall also increase. These two effects shall balance each other and the benefits acquired from the bond would remain the same in the long run, due to the pull-to-par effect of the bond.